OpenStudy (anonymous):

Please help check my answer: describe the line passing through (-5,0,4) and (6,-3,2):: => l(t) = (-5+11t)i - (3t)j + (4-2t)k

OpenStudy (phi):

yes. the direction vector is pt1 - pt2 (or vice versa) (6,-3,2) - (-5, 0, 4) = (11, -3, -2) pick either point as the starting point: P = (-5,0,4)+ t(11,-3,-2) which is your version. we could also do P= (6,-3,2)+ t(11,-3,-2) or P= (6,-3,2)+ t(-11,3,2) (changed the signs of the direction vector)